OpenStudy (anonymous):

Solve x^2 + 8x - 48 = 0 by completing the square

OpenStudy (anonymous):

the answer equals to (x-12)(x+4)=0 then we have x-12=0 and x+4=0, then we have two correct answers x=12 and x=(-4).

OpenStudy (anonymous):

so is this right ? x2 + 8x - 48 = 0 x2 + 8x = 48 x² + 8x + ? = 48 + ? The object is to force the left side of the equation to be a perfect square trinomial. Take half of the coefficient of the linear x term (it is 8) --> 4 Square it. 4² = 16 x² + 8x + 16 = 48 + 16 (x + 4)^2= 64 (x + 4) = + or - radical 64 (x+4) = + or - 8 (x-12) (x+4) = 0

OpenStudy (anonymous):

yep that's it.

OpenStudy (anonymous):

=>X^2+8x-48=0 =>x^2+8x=48 =>x^2+8x+16=48+16 =>(x+4)^2=64 =>x+4=+-\/64 =>x=4+-8 =>x=-4+8 or x=-4-8